OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (-3, -5) and (-11, -7)?

OpenStudy (rmrjr22):

(x-a)^2 + (y-b)^2 = r^2

OpenStudy (rmrjr22):

(x2+x1)/2,(y2+y1)/2=midpoint

OpenStudy (rmrjr22):

plug and play into midpoint formula

OpenStudy (anonymous):

idk what that means haha

OpenStudy (rmrjr22):

(-3,5) = (x1,y1) (-11,-7) = (x2,y2)

hero (hero):

If you don't know what it means, use graph paper.

OpenStudy (rmrjr22):

in the formula you are taking the difference of two points then dividing it by two to get the center of those two points you just subtracted from each other...

OpenStudy (anonymous):

-14,2?

OpenStudy (rmrjr22):

-7+ 5 = ?

OpenStudy (rmrjr22):

-7 + (-5) = ? *********

OpenStudy (anonymous):

-2

OpenStudy (rmrjr22):

-12

OpenStudy (rmrjr22):

sorry my fault

hero (hero):

If you use graph paper, you'll get something like this: 

hero (hero):

The line segment will intersect a point that is the midpoint of the segment.

OpenStudy (anonymous):

-7,-6?

hero (hero):

Bingo

OpenStudy (rmrjr22):

Bingo!

hero (hero):

Bingo is one of my favorite games

OpenStudy (rmrjr22):

it will be mine when i get older

hero (hero):

You assume only old people play it.

OpenStudy (anonymous):

bingo is fun